OpenStudy (anonymous):

I have a quadratic eq. question: Which of the following is a solution of 2x2 = -3x - 5? a. x = -3 -7i over 4 b. x = 3 + 7i over 4 c. x = 3 - i sqrt 31 over 4 d. x = -3 + i sqrt 31 over 4

OpenStudy (anonymous):

here agin a=+2 b=+3 and c=+5

OpenStudy (anonymous):

check if you get d. x = -3 + i sqrt 31 over 4 after using quadratic formula..

OpenStudy (anonymous):

so it got c. but it's because of the negative sign again

OpenStudy (anonymous):

check if you get d. x = -3 + i sqrt 31 over 4 after using quadratic formula

OpenStudy (anonymous):

check the sign of b

OpenStudy (anonymous):

you have to make the 3x and the 5 positive right!

OpenStudy (anonymous):

no not that..

OpenStudy (anonymous):

to get it in the form of the quadratic equation? with zero on the other side?

OpenStudy (anonymous):

yup 2x2 = -3x - 5 2x2 + 3x + 5 =0 now you compare it with standard quadratic..

OpenStudy (anonymous):

ok yes it get this now! you're a great teacher! then it's just factoring it out. the sqrt's upset me lol!

OpenStudy (anonymous):

sometimes it can be factored other times not..

OpenStudy (anonymous):

so you just simplify as much as possible and that's your answer?

OpenStudy (anonymous):

thanks again I am beginning to understand..

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

but for ax^2+bx+c = 0 x= (-b - sqrt(b^2 -4ac))/(2a) , (-b + sqrt(b^2 -4ac))/(2a)

OpenStudy (anonymous):

oh cool, yes I can see now! I am putting that in my notes now. You are super!

OpenStudy (anonymous):

yw